# VAEmo: Efficient Representation Learning for Visual-Audio Emotion with Knowledge Injection

### v1

@@ Abstract.
Audiovisual emotion recognition (AVER) aims to infer human emotions from nonverbal visual-audio (VA) cues, offering modality-complementary and language-agnostic advantages. However, AVER remains challenging due to the inherent ambiguity of emotional expressions, cross-modal expressive disparities, and the scarcity of reliably annotated data. Recent self-supervised AVER approaches have introduced strong multimodal representations, yet they predominantly rely on modality-specific encoders and coarse content-level alignment, limiting fine-grained emotional semantic modeling. To address these issues, we propose VAEmo, an efficient two-stage framework for emotion-centric joint VA representation learning with external knowledge injection. In Stage 1, a unified and lightweight representation network is pre-trained on large-scale speaker-centric VA corpora via masked reconstruction and contrastive objectives, mitigating the modality gap and learning expressive, complementary representations without emotion labels. In Stage 2, multimodal large language models automatically generate detailed affective descriptions according to our well-designed chain-of-thought prompting for only a small subset of VA samples; these rich textual semantics are then injected by aligning their corresponding embeddings with VA representations through dual-path contrastive learning, further bridging the emotion gap. Extensive experiments on multiple downstream AVER benchmarks show that VAEmo achieves state-of-the-art performance with a compact design, highlighting the benefit of unified cross-modal encoding and emotion-aware semantic guidance for efficient, generalizable VA emotion representations. Source code and pre-trained models will be available at https://github.com/MSA-LMC/VAEmo
@@ 1. Introduction
To overcome these limitations in emotion recognition, this field has increasingly turned to self-supervised learning (SSL) techniques (He et al., 2022; Tong et al., 2022; Huang et al., 2022; Radford et al., 2021; Hsu et al., 2021). By leveraging large amounts of unlabeled audio-visual data, SSL approaches have shown significant potential for learning highly generalizable multimodal representations (Gong et al., [n. d.]; Huang et al., 2023; Zellers et al., 2021), proficient at capturing general audio-visual correspondences. However, they overlook the subtle emotional cues, resulting in mismatch to emotion recognition tasks, as these approaches lack a specific design of emotion-specific feature extraction and fine-grained emotional semantic alignment during pre-training. Consequently, models derived from these self-supervised manner often miss emotional nuances, unable to effectively capture intra-modal emotional cues or align fine-grained emotional semantics across modalities in AVER downstream tasks.
@@
Our knowledge injection pipeline leverages specialized MLLMs to generate emotion-focused descriptions from both modalities, creating a rich semantic bridge between raw signals and emotional concepts. Through a dual-path contrastive learning approach that maintains separate audio-text and video-text learning paths while optimizing the same representation network, we preserve modality-specific characteristics while enriching emotional semantics. Our carefully designed two-stage training paradigm effectively addresses the gap between general audio-visual representation learning and emotion-specific feature extraction. Extensive experiments demonstrate that VAEmo consistently outperforms previous state-of-the-art approaches while using significantly fewer parameters. In summary, our work makes the following contributions:
@@
An innovative knowledge injection method leveraging MLLMs to generate fine-grained affective descriptions , creating a rich semantic bridge between raw signals and emotional concepts through a dual-path contrastive learning approach.
Comprehensive experimental validation across 6 AVER datasets, architectural simplicity coupled with targeted knowledge injection yields superior results compared to more complex models.
@@ 2. Related Work
Audio-Visual Emotion Recognition. Audio-visual emotion recognition (AVER) integrates speech and facial expression analysis, providing complementary multi-modal perspectives on emotional understanding (Mehrabian, 2017; Wu et al., 2014). Early approaches relied on supervised learning with hand-crafted features, such as eGeMAPS (Eyben et al., 2015) for audio and VGGFace (Cao et al., 2018) for visual features. With the advancement of deep learning, end-to-end approaches became prevalent. Liu et al. (Liu et al., 2022) proposed T-ESFL, which employs collaborative multi-task learning to improve AVER performance. Zhang et al. (Zhang et al., 2023b) developed T-MEP, introducing attention mechanisms to enhance cross-modal interaction. Recently, self-supervised approaches have made significant breakthroughs, such as AW-HuBERT (Tran et al., 2023) leveraged audio-visual self-supervised pre-training to enhance scene adaptation capabilities, and HiCMAE (Sun et al., 2024a) learned expressive representations through hierarchical contrastive masked autoencoding. Additionally, large models have been applied to emotion analysis. Chen et al. (Chen et al., 2024a) introduced FineCLIPER, which utilizes LLMs to generate emotional annotations, and Zhao et al. (Zhao and Patras, 2023) proposed DFER-CLIP, which explores visual prompt tuning for emotion recognition. In contrast to these approaches, our VAEmo adopts a compact single-stream architecture and a two-stage training paradigm, significantly enhancing emotional representation capacity while maintaining high efficiency.
@@
Self-Supervised Learning and Knowledge Injection. Self-supervised learning (SSL) has become the dominant paradigm for representation learning. In the visual domain, MAE (He et al., 2022) employs high-ratio masked reconstruction pre-training, while Tong et al. (Tong et al., 2022) extended this to the spatiotemporal domain. For audio, Hsu et al.’s (Hsu et al., 2021) HuBERT and Chen et al.’s (Chen et al., 2022) WavLM utilize hidden unit prediction to establish powerful speech representations. In multi-modal learning, Gong et al. (Gong et al., [n. d.]) combined contrastive learning with MAE, while MAViL (Huang et al., 2023) explored more on contrastive pre-training. Despite these advances, existing SSL approaches primarily focus on general-purpose representation learning, often overlooking the injection of domain-specific knowledge such as emotion semantics. For knowledge injection, with the emergence of large models, researchers have begun exploring LLMs as external knowledge sources. LLM2CLIP (Wu et al., [n. d.]) injects knowledge from large models into CLIP (Radford et al., 2021) models to enhance the performance of visual-language models. MATE (Jang et al., 2024) uses LoRA-tuned (Shen et al., [n. d.]) large language models to enhance the performance of CLIP visual encoders.
Different from these works, our approach uniquely leverages multimodal LLMs (Qwen2-VL (Wang et al., 2024) and Qwen2-Audio (Chu et al., 2024)) to generate emotion-focused descriptions, incorporating emotional knowledge through dual-path contrastive learning that effectively bridges the semantic gap between general and emotion-specific representations.
@@ 3. Approach
Overview. As shown in Fig. 2, our approach follows a two-stage training paradigm. Raw inputs are encoded through our representation network, which consists of three key components: modality-specific tokenizers, a feature transformation encoder, and a cross-modal fusion encoder. Based on different training stages, the encoded audio-visual features are input to stage-specific task heads. In Stage 1, this includes a combination of masked autoencoder decoder and contrastive learning objectives, facilitating the establishment of cross-modal representations. In Stage 2, this becomes our knowledge injection framework, where captions generated by our carefully designed Chain of Thought (CoT)-based caption generation mechanism serve as emotional knowledge injected into the representation network. The following sections provide detailed descriptions of VAEmo’s architecture and our two-stage training methodology.
@@ 3.1. Network Architecture
Our VAEmo architecture consists of four main components as illustrated in Fig. 2: (1) Modality-specific tokenizers : convert raw audio and visual inputs into token sequences with consistent dimensions; (2) A shared-parameter feature transformation encoder : processes each modality while sharing the same parameters across modalities; (3) A cross-modal fusion encoder : integrates information by concatenating outputs from across modalities, also utilizing shared parameters; and (4) Stage-specific heads : vary depending on the training stage .
Both the feature transformation encoder and fusion encoder are implemented using a Vision Transformer (ViT)-based architecture. This intentionally simplified design offers two significant advantages: (1) it allows us to focus research attention on our novel two-stage training methodology rather than architectural innovations, providing clearer insights into the effectiveness of our knowledge injection approach; and (2) the unified architecture with shared parameters across modalities enables tighter integration of multimodal features in our encoder, while drastically reducing the model’s parameter count compared to traditional dual-stream approaches commonly used in audio-visual representation learning.
@@
Our model architecture consists of a unified representation network and stage-specific task heads that operate on multimodal inputs:
where and represent the modality-specific tokenizers, represents the representation network parameters, and denotes token-wise concatenation of features. The complete model for a specific training stage can be expressed as:
@@
Stage 1: A dual-purpose head () supporting both masked reconstruction and contrastive learning objectives. The reconstruction component decodes masked tokens while the contrastive component aligns audio-visual features
@@ 3.2. Training Paradigm
Stage 1: Contrastive Learning and Masked Self-Supervised Learning. For Stage 1 pre-training, we employ a dual strategy combining contrastive learning and masked self-supervised learning, an approach proven effective in prior works like CAV-MAE (Gong et al., [n. d.]) and MAViL (Huang et al., 2023). Our contrastive objective aligns global representations from audio and video modalities using standard InfoNCE loss with temperature scaling. Simultaneously, we perform masked self-supervised learning by randomly masking tokens (80% for audio, 90% for video) and tasking the model to reconstruct the original tokens from these partially masked inputs. The reconstruction process involves processing masked inputs through our encoders, fusing modality-specific features, and computing the mean squared error between original and reconstructed tokens. This combined approach offers complementary benefits: contrastive learning captures high-level cross-modal correspondences, while reconstruction preserves fine-grained details within each modality. After Stage 1 pre-training, we discard the reconstruction decoder and retain only the representation network for Stage 2, allowing the model to first learn generalizable speech-face representations before incorporating emotional knowledge.
@@
Stage 2: Knowledge Injection with MLLMs. In the second stage, we further enhance our model’s emotion representation capabilities through knowledge injection. We develop a fully automated knowledge extraction and injection pipeline leveraging Multimodal Large Language Models (MLLMs). Using modality-specific MLLMs—Qwen2-VL for video and Qwen2-Audio for audio—we generate rich emotional descriptions through a carefully designed Chain-of-Thought (CoT) prompting strategy. Our CoT process first guides MLLMs to describe observable physical characteristics—facial action units, overall posture, and body movements for video, or acoustic properties for audio. Based on these objective features, MLLMs then assess emotional states, including intensity (arousal) and valence, finally analyzing overall emotional changes throughout the sequence. To mitigate potential hallucinations, we explicitly instruct models to ground inferences in observable features, perform multiple inference passes with voting mechanisms for consensus on emotional states, and employ a secondary LLM to extract and filter emotion-irrelevant information. This comprehensive approach ensures generated descriptions remain faithful to inputs while focusing on affective content, producing specialized affective vocabulary (e.g., ”assertiveness”, ”amusement”, ”demeanor”) that captures subtle emotional nuances and complex compound emotions, as illustrated in Fig. 3.
@@
Since the generated captions are much longer than CLIP (Radford et al., 2021)’s 77-token limit (in our implementation, tokens counted using GPT-2 (Radford et al., [n. d.]) tokenizer can reach up to 100) and contain specialized emotional vocabulary absent from CLIP’s training data, we employ LLM-based embeddings for text representation. Specifically, we use the NV-Embed-v2 (Lee et al., 2024) model, which adapts a decoder-only LLM as an embedding generator by employing specialized latent attention layers and removing causal attention masks. Our ablation experiments demonstrate that LLM embeddings significantly outperform both CLIP text embeddings and BERT embeddings for our emotion recognition tasks.
After encoding these captions into text embeddings and for audio and video, respectively, we implement a dual-path contrastive learning caption-audio/video approach (Fig. 2) that replaces the traditional audio-visual contrastive objective from Stage 1. This strategy creates two parallel contrastive objectives—audio-text and video-text pairs:
@@
A crucial aspect of this dual-path approach is that both objectives simultaneously optimize the same representation network through different gradient paths. Since both and are outputs from the representation network, the gradients from the two contrastive losses flow back to update the same set of parameters. This creates a rich training signal that enhances the encoder’s ability to capture emotional features across modalities while preserving their unique characteristics.
This shared parameter optimization is particularly effective for emotion recognition for several reasons: 1. The encoder learns to project both modalities into a common emotional semantic space defined by the MLLM-generated captions; 2. The parallel gradient flows provide complementary supervision that enhances the robustness of learned representations; 3. The shared parameters facilitate cross-modal transfer, allowing emotional patterns learned from one modality to benefit the other.
After completing both training stages, we extract the encoder and fine-tune it on downstream tasks to leverage the emotion-rich representations for specific applications. The sequential nature of our training paradigm—from general audio-visual correspondences to emotion-specific knowledge—creates a robust foundation for emotion recognition that outperforms more complex architectural alternatives, as demonstrated in our experimental results.
@@ 4. Experiments
In this section, we evaluate the effectiveness of VAEmo on a suite of audio-visual emotion recognition (AVER) tasks. Our experiments cover both categorical and dimensional emotion recognition paradigms across diverse datasets, spanning from controlled lab environments to unconstrained in-the-wild scenarios. This comprehensive evaluation framework is designed to rigorously assess the generalization ability of the learned emotion representations.

### v2

@@ Abstract.
Audiovisual emotion recognition (AVER) aims to infer human emotions from nonverbal visual-audio (VA) cues, offering modality-complementary and language-agnostic advantages. However, AVER remains challenging due to the inherent ambiguity of emotional expressions, cross-modal expressive disparities, and the scarcity of reliably annotated data. Recent self-supervised AVER approaches have introduced strong multimodal representations, yet they predominantly rely on modality-specific encoders and coarse content-level alignment, limiting fine-grained emotional semantic modeling. To address these issues, we propose VAEmo, an efficient two-stage framework for emotion-centric joint VA representation learning with external knowledge injection. In Stage 1, a unified and lightweight representation network is pre-trained on large-scale speaker-centric VA corpora via masked reconstruction and contrastive objectives, mitigating the modality gap and learning expressive, complementary representations without emotion labels. In Stage 2, multimodal large language models automatically generate detailed affective descriptions according to our well-designed chain-of-thought prompting for only a small subset of VA samples; these rich textual semantics are then injected by aligning their corresponding embeddings with VA representations through dual-path contrastive learning, further bridging the emotion gap. Extensive experiments on multiple downstream AVER benchmarks show that VAEmo achieves state-of-the-art performance with a compact design, highlighting the benefit of unified cross-modal encoding and emotion-aware semantic guidance for efficient, generalizable VA emotion representations. Source code and pre-trained models will be available at https://github.com/MSA-LMC/VAEmo.
@@ 1. Introduction
To overcome these limitations in emotion recognition, this field has increasingly turned to self-supervised learning (SSL) techniques (He et al., 2022; Tong et al., 2022; Huang et al., 2022; Radford et al., 2021; Hsu et al., 2021). By leveraging large amounts of unlabeled audio-visual data, SSL methods have shown great potential for learning rich multimodal representations without requiring expensive emotion annotations (Gong et al., [n. d.]; Huang et al., 2023; Zellers et al., 2021). However, existing SSL methods, while proficient at capturing general audio-visual correspondences, often overlook subtle emotional cues that require specialized attention. This creates a fundamental mismatch for emotion recognition, as these methods fail to target the extraction of emotion-specific features and fine-grained emotional semantic alignment during pre-training. Consequently, models derived from these self-supervised manner often miss emotional nuances, unable to effectively capture intra-modal emotional cues or align fine-grained emotional semantics across modalities in AVER downstream tasks.
@@
Our knowledge injection pipeline leverages specialized MLLMs to generate emotion-focused descriptions from both modalities, creating a rich semantic bridge between raw signals and emotional concepts. Through a dual-path contrastive learning approach that maintains separate audio-text and video-text learning paths while optimizing the same representation network, we preserve modality-specific characteristics while enriching emotional semantics. Our carefully designed two-stage training paradigm effectively addresses the gap between general audio-visual representation learning and emotion-specific feature extraction. By first establishing strong cross-modal representations in Stage 1 before injecting emotion-specific knowledge in Stage 2, we create a more stable learning trajectory. Extensive experiments demonstrate that VAEmo consistently outperforms previous state-of-the-art approaches while using significantly fewer parameters. In summary, our work makes the following contributions:
@@
An innovative knowledge injection method leveraging MLLMs to generate emotion-focused descriptions, creating a rich semantic bridge between raw signals and emotional concepts through a dual-path contrastive learning approach.
Comprehensive experimental validation revealing that for emotion-specific tasks, architectural simplicity coupled with targeted knowledge injection yields superior results compared to more complex models.
@@ 2. Related Work
Audio-Visual Emotion Recognition. Audio-visual emotion recognition (AVER) integrates speech and facial expression analysis, providing complementary multi-modal perspectives on emotion understanding (Mehrabian, 2017; Wu et al., 2014). Early approaches relied on supervised learning with hand-crafted features, such as eGeMAPS (Eyben et al., 2015) for audio and VGGFace (Cao et al., 2018) for visual features. With the advancement of deep learning, end-to-end approaches became prevalent. Liu et al. (Liu et al., 2022) proposed T-ESFL, which employs collaborative multi-task learning to improve AVER performance. Zhang et al. (Zhang et al., 2023b) developed T-MEP, introducing attention mechanisms to enhance cross-modal interaction. Recently, self-supervised approaches have made significant breakthroughs, such as AW-HuBERT (Tran et al., 2023) leveraged audio-visual self-supervised pre-training to enhance scene adaptation capabilities, and HiCMAE (Sun et al., 2024a) learned expressive representations through hierarchical contrastive masked autoencoding. Additionally, large models have been applied to emotion analysis. Chen et al. (Chen et al., 2024a) introduced FineCLIPER, which utilizes LLMs to generate emotional annotations, and Zhao et al. (Zhao and Patras, 2023) proposed DFER-CLIP, which explores visual prompt tuning for emotion recognition. In contrast to these methods, our VAEmo adopts a compact single-stream architecture and two-stage training paradigm, significantly enhancing emotional representation capacity while maintaining high efficiency.
@@
Self-Supervised Learning and Knowledge Injection. Self-supervised learning (SSL) has become the dominant paradigm for representation learning. In the visual domain, MAE (He et al., 2022) employs high-ratio masked reconstruction pre-training, while Tong et al. (Tong et al., 2022) extended this to the spatiotemporal domain. For audio, Hsu et al.’s (Hsu et al., 2021) HuBERT and Chen et al.’s (Chen et al., 2022) WavLM utilize hidden unit prediction to establish powerful speech representations. In multi-modal learning, Gong et al. (Gong et al., [n. d.]) combined contrastive learning with MAE, while MAViL (Huang et al., 2023) explored more on contrastive object. Despite these advances, existing SSL approaches primarily focus on general-purpose representation learning, often overlooking the injection of domain-specific knowledge such as emotion semantics. For knowledge injection, with the emergence of large models, researchers have begun exploring LLMs as external knowledge sources. LLM2CLIP (Wu et al., [n. d.]) injects knowledge from large models into CLIP (Radford et al., 2021) models to enhance the performance of visual-language models. MATE (Jang et al., 2024) uses LoRA-tuned (Shen et al., [n. d.]) large language models to enhance the performance of CLIP visual encoders. Different from these works, our method uniquely uses multimodal LLMs (Qwen2-VL (Wang et al., 2024) and Qwen2-Audio (Chu et al., 2024)) to generate emotion-focused descriptions, incorporating emotional knowledge through dual-path contrastive learning that effectively bridges semantic gap between general and emotion-specific representations.
@@ 3. Approach
Overview. As shown in Fig. 2, our method follows a two-stage training paradigm. Raw inputs are encoded through our representation network, which consists of three key components: modality-specific tokenizers, a feature transformation encoder, and a cross-modal fusion encoder. Based on different training stages, the encoded audio and visual features are input to stage-specific task heads. In Stage 1, this includes a combination of masked reconstruction and contrastive learning objectives, facilitating the establishment of cross-modal representations. In Stage 2, this becomes our knowledge injection framework, where captions generated by our carefully designed Chain of Thought (CoT)-based caption generation mechanism serve as emotional knowledge injected into representation network. The following sections provide detailed descriptions of VAEmo’s architecture and two-stage training paradigm.
@@ 3.1. Network Architecture
VAEmo architecture consists of four main components as shown in Fig. 2: (1) Modality-specific tokenizers : convert raw audio and visual inputs into token sequences with consistent dimensions; (2) A shared-parameter feature transformation encoder : encode each modality with multiple Transformer layers using shared parameters. (3) A cross-modal fusion encoder : integrates information by concatenating outputs from across modalities, followed by Transformer layers with shared parameters; (4) Stage-specific heads : vary depending on the training stage .
Both the feature transformation encoder and fusion encoder are implemented using a Vision Transformer (ViT)-based architecture. This intentionally simplified design offers two significant advantages: (1) it allows us to focus research attention on our novel two-stage training methodology rather than architectural innovations, providing clearer insights into the effectiveness of our knowledge injection approach; and (2) the unified architecture with shared parameters across modalities enables tighter integration of multimodal features in our encoder, while reducing the model’s parameter count compared to traditional dual-stream methods commonly used in audio-visual representation learning.
@@
Our model architecture consists of a unified representation network and stage-specific task heads .
where and represent the modality-specific tokenizers, represents the representation network parameters, and denotes token-wise concatenation of feature sequences. The data flow for a specific training stage can be expressed as:
@@
Stage 1: A dual-purpose head () supporting both masked reconstruction and contrastive learning objectives. The reconstruction component decodes masked tokens while the contrastive component aligns audio-visual representation.
@@ 3.2. Training Paradigm
Stage 1: Contrastive Learning and Masked Reconstruction. For Stage 1 pre-training, we employ a dual strategy combining contrastive learning and masked self-supervised learning, an approach proven effective in prior works like CAV-MAE (Gong et al., [n. d.]) and MAViL (Huang et al., 2023). Our contrastive objective employs the standard InfoNCE loss with temperature scaling to align the global representations of paired video and audio. For each video–audio pair, the global representations for video and audio are obtained by token-wise mean pooling over the output feature sequences and , respectively. Simultaneously, we perform masked self-supervised learning by randomly masking tokens (80% for audio, 90% for video) and tasking the model to reconstruct the original data from these partially masked inputs. The reconstruction process involves processing masked inputs through our encoders, fusing modality-specific features, and computing the mean squared error loss between original and reconstructed data. This combined approach offers complementary benefits: contrastive learning captures high-level cross-modal correspondences, while reconstruction preserves fine-grained details within each modality. After Stage 1 pre-training, we discard the reconstruction decoder and retain only the representation network for Stage 2, allowing the model to first learn generalizable speech-face representations before incorporating emotional knowledge.
Stage 2: Knowledge Injection with MLLMs. In the second stage, we further enhance our model’s emotion representation capabilities based on the stage 1 parameters through knowledge injection. We develop a fully automated knowledge extraction and injection pipeline leveraging Multimodal Large Language Models (MLLMs). Using modality-specific MLLMs—Qwen2-VL for video and Qwen2-Audio for audio, we generate rich emotional descriptions through a carefully designed Chain-of-Thought (CoT) prompting strategy.
@@
Our CoT process first guides MLLMs to describe observable physical characteristics—facial action units, overall posture, and body movements for video, or acoustic properties for audio. Based on these objective features, MLLMs then assess emotional states, including intensity (arousal) and valence, and finally analyze overall emotional changes throughout the sequence. To ensure that the generated captions are both factual and emotionally relevant, we introduce a refinement mechanism. First, we explicitly instruct the MLLMs to ground all inferences in observable features. Then, we conduct multiple inference passes per input and apply a majority voting strategy to filter out inconsistent or hallucinated outputs. We employ a secondary LLM (GPT-4o-mini (Hurst et al., 2024)) to filter the results by discarding emotion-irrelevant content.This refinement pipeline encourages the generation of expressive and specialized affective vocabulary (e.g., “assertiveness,” “amusement,” “demeanor”), enabling the model to capture subtle and compound emotional states. As illustrated in Fig. 3, these high-quality descriptions serve as valuable supervision signals during training.
@@
Since the generated captions are much longer than CLIP (Radford et al., 2021)’s 77-token limit (in our implementation, tokens counted using GPT-2 (Radford et al., [n. d.]) tokenizer can reach up to 100) and contain specialized emotional vocabulary absent from CLIP’s training data, we employ LLM-based embeddings for text representation. Specifically, we use the NV-Embed-v2 (Lee et al., 2024) model, which adapts a decoder-only LLM as an embedding generator by employing specialized latent attention layers and removing causal attention masks. Our ablation experiments demonstrate that LLM embeddings significantly outperform both CLIP text embeddings and BERT embeddings for our emotion recognition tasks. After encoding the captions into text embeddings, we further process each embedding through a modality-specific adapter module, resulting in and for audio and video, respectively. The adapter design follows the implementation in LLM2CLIP (Wu et al., [n. d.]). We then implement a dual-path contrastive learning approach (Fig. 2), which replaces the traditional audio-visual contrastive objective from Stage 1. Specifically, we construct two parallel contrastive objectives: one between audio caption embedding and audio global representation, and the other between video caption embedding and video global representation.
@@
A crucial aspect of this dual-path approach is that both objectives simultaneously optimize the same encoder through different gradient paths. and are the token-wise mean-pooled embedding of the feature transformation encoder output for audio and video, respectively. The gradients from the two contrastive losses flow back to update the same set of parameters. This creates a rich training signal that enhances the encoder’s ability to aware emotional knowledge for both modalities.
After completing both training stages, we extract the encoder and fine-tune it on downstream tasks to leverage the emotion-rich representations for specific applications. The sequential nature of our training paradigm—from general audio-visual correspondences to emotion-specific knowledge—creates a robust foundation for visual-audio emotion recognition.
@@ 4. Experiments
In this section, we evaluate the effectiveness of VAEmo on a suite of audio-visual emotion recognition (AVER) tasks. Our experiments cover both categorical and dimensional emotion recognition paradigms across diverse datasets, spanning from controlled lab environments to unconstrained in-the-wild scenarios.